OpenStudy (anonymous):

What are the exact solution of x^2=5x+2?

OpenStudy (mathteacher1729):

How far did you get with this problem before getting stuck?

OpenStudy (anonymous):

no where

OpenStudy (anonymous):

x^2=5x+2; 0 = X^2-5*X-2 use the Quadratic formula: X = (-b ±√(b²-4ac))/(2a) = (5±√(25 +8))/(2) positive discriminant = 33; two real roots: X=-0.3722813232690143 and 5.372281323269014

OpenStudy (mathteacher1729):

If you go to http://www.wolframalpha.com and type solve x^2=5x+2 then click "show steps" it will solve it and show you the steps.